# The Mixed Integer Trust Region Problem

### v1

@@ 7.2 Approximation algorithm for Equation MIQP
and note that $\mathcal{S}$ constitutes the feasible region of Equation MIQP. We apply theorem 1 in [6]. If we find out that $\mathcal{S}$ is empty, we are done. Otherwise, we obtain $p^{\prime}\in\left\{0,1,\dots,p\right\}$, $n^{\prime}\in\left\{p^{\prime},p^{\prime}+1,\dots,p^{\prime}+n-p\right\}$, a map $\tau:\mathbb{R}^{n^{\prime}}\to\mathbb{R}^{n}$ of the form $\tau(x^{\prime})=\bar{x}+Mx^{\prime}$, with $\bar{x}\in\mathbb{Z}^{p}\times\mathbb{Q}^{n-p}$ and $M\in\mathbb{Q}^{n\times n^{\prime}}$ of full rank, such that the polyhedron
@@
We perform the change of variables $x=\bar{x}+Mx^{\prime}$ to Equation MIQP and, after dropping the constant $\bar{x}^{\mathsf{T}}H\bar{x}+h^{\mathsf{T}}\bar{x}$ in the objective function, we obtain the problem
$\begin{split}\min&\quad{x^{\prime}}^{\mathsf{T}}H^{\prime}x^{\prime}+{h^{\prime}}^{\mathsf{T}}x^{\prime}\\ \operatorname{s.t.}&\quad W^{\prime}x^{\prime}\leq w^{\prime}\\ &\quad x^{\prime}\in\mathbb{Z}^{p^{\prime}}\times\mathbb{R}^{n^{\prime}-p^{\prime}},\end{split}$ (${}^{\prime}$)
where $H^{\prime}:=M^{\mathsf{T}}HM$, ${h^{\prime}}^{\mathsf{T}}:=h^{\mathsf{T}}M+2\bar{x}^{\mathsf{T}}HM$, $W^{\prime}:=WM$, and $w^{\prime}:=w-W\bar{x}$. Since $\mathcal{P}$ is bounded, also the polyhedron $\mathcal{P}^{\prime}$ is bounded. From theorem 1 in [6], the polyhedron $\mathcal{P}^{\prime}$ is full-dimensional.
Since the definition of $\epsilon$-approximate solution is preserved under changes of variables and translations of the objective function, it then suffices to find an $\epsilon$-approximate solution to Equation ${}^{\prime}$. For ease of notation, from now on we consider Equation MIQP instead of Equation ${}^{\prime}$ and we further assume, without loss of generality, that $\{x\in\mathbb{R}^{n}:Wx\leq w\}$ in Equation MIQP is full-dimensional.
@@
$\left\{d^{\mathsf{T}}x:Wx\leq w\right\}\in\left[\rho,\rho+20(n+1)p2^{p(p-1)/4}\right].$
@@
$\{x\in\mathbb{R}^{n}:{d}^{\mathsf{T}}x=\beta\},\qquad\beta=\lceil\rho\rceil,\ \lceil\rho\rceil+1,\dots,\lceil\rho+20(n+1)p2^{p(p-1)/4}\rceil.$
For each $\beta=\lceil\rho\rceil,\dots,\lceil\rho+20(n+1)p2^{p(p-1)/4}\rceil$, we define the optimization problem
@@
For each $\beta$, we apply theorem 1 in [6] to the feasible region of Equation MIQP-$\beta$, as in Step 2. If we find out that the feasible region is empty, we are done. Otherwise, we transform, with a change of variables, Equation MIQP-$\beta$ into a new problem, which we denote by Problem MIQP-$\beta^{\prime}$. Since $d_{p+1}=\cdots=d_{n}=0$, Problem MIQP-$\beta^{\prime}$ has at most $p-1$ integer variables and $n-p$ continuous variables.
Step 3: Recursion. We apply recursively the algorithm, from Step 1, to each obtained Problem MIQP-$\beta^{\prime}$, for each $\beta=\lceil\rho\rceil,\dots,\lceil\rho+20(n+1)p2^{p(p-1)/4}\rceil$. Since each time the problem is partitioned in Step 2, the number of integer variables decreases by one, the total number of iterations of the algorithm is upper bounded by
$\left(20(n+1)p2^{p(p-1)/4}\right)^{p}.$
Step 4: Termination. Among all the feasible solutions to Equation MIQP found, we return one with the minimum objective value. The result then follows from Lemma 15. This is because the function $f(x):=x^{\mathsf{T}}Hx+h^{\mathsf{T}}x$, from the original feasible region $\{x\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}:Wx\leq w\}$ to $\mathbb{R}$, is lower bounded by assumption, and thus has a (global) minimum, due to theorem 4 in [7]. ∎

### v2

@@ 8.2 Approximation algorithm for Equation MIQP
and note that $\mathcal{S}$ constitutes the feasible region of Equation MIQP. We apply the algorithm in theorem 1 in [8]. The algorithm either returns that $\mathcal{S}$ is empty, or it returns $p^{\prime}\in\left\{0,1,\dots,p\right\}$, $n^{\prime}\in\left\{p^{\prime},p^{\prime}+1,\dots,p^{\prime}+n-p\right\}$, a map $\tau:\mathbb{R}^{n^{\prime}}\to\mathbb{R}^{n}$ of the form $\tau(x^{\prime})=\bar{x}+Mx^{\prime}$, with $\bar{x}\in\mathbb{Z}^{p}\times\mathbb{Q}^{n-p}$ and $M\in\mathbb{Q}^{n\times n^{\prime}}$ of full rank, such that the polyhedron
@@
If $\mathcal{S}$ is empty we are done. Otherwise, we perform the change of variables $x=\bar{x}+Mx^{\prime}$ to Equation MIQP and, after dropping the constant $\bar{x}^{\mathsf{T}}H\bar{x}+h^{\mathsf{T}}\bar{x}$ in the objective function, we obtain the problem
$\begin{split}\min&\quad{x^{\prime}}^{\mathsf{T}}H^{\prime}x^{\prime}+{h^{\prime}}^{\mathsf{T}}x^{\prime}\\ \operatorname{s.t.}&\quad W^{\prime}x^{\prime}\leq w^{\prime}\\ &\quad x^{\prime}\in\mathbb{Z}^{p^{\prime}}\times\mathbb{R}^{n^{\prime}-p^{\prime}},\end{split}$ (MIQP)
where $H^{\prime}:=M^{\mathsf{T}}HM$, ${h^{\prime}}^{\mathsf{T}}:=h^{\mathsf{T}}M+2\bar{x}^{\mathsf{T}}HM$, $W^{\prime}:=WM$, and $w^{\prime}:=w-W\bar{x}$. Since $\mathcal{P}$ is bounded, also the polyhedron $\mathcal{P}^{\prime}$, which is full-dimensional, is also bounded.
Since the definition of $\epsilon$-approximate solution is preserved under changes of variables and translations of the objective function, it then suffices to find an $\epsilon$-approximate solution to Equation MIQP. For ease of notation, from now on we consider Equation MIQP instead of Equation MIQP and we further assume, without loss of generality, that $\{x\in\mathbb{R}^{n}:Wx\leq w\}$ in Equation MIQP is full-dimensional.
@@
$\left\{d^{\mathsf{T}}x:Wx\leq w\right\}\subseteq\left[\rho,\rho+14(n+1)p2^{p(p-1)/4}\right].$
@@
$\{x\in\mathbb{R}^{n}:{d}^{\mathsf{T}}x=\beta\},\qquad\beta=\lceil\rho\rceil,\ \lceil\rho\rceil+1,\dots,\lfloor\rho+14(n+1)p2^{p(p-1)/4}\rfloor.$
For each $\beta=\lceil\rho\rceil,\lceil\rho\rceil+1,\dots,\lfloor\rho+14(n+1)p2^{p(p-1)/4}\rfloor$, we define the optimization problem
@@
If $n=1$, Equation MIQP-$\beta$ can be trivially solved, thus we now assume $n\geq 2$. For each $\beta$, we apply theorem 1 in [8] to the feasible region of Equation MIQP-$\beta$. If we find out that the feasible region is empty, we are done. Otherwise, we transform, with a change of variables, Equation MIQP-$\beta$ into a new problem, which we denote by Problem MIQP-$\beta^{\prime}$. Since $d_{p+1}=\cdots=d_{n}=0$, Problem MIQP-$\beta^{\prime}$ has at most $p-1$ integer variables and $n-p$ continuous variables.
Step 3: Recursion. We apply recursively the algorithm, from Step 1, to each obtained Problem MIQP-$\beta^{\prime}$, for each $\beta=\lceil\rho\rceil,\lceil\rho\rceil+1,\dots,\lfloor\rho+14(n+1)p2^{p(p-1)/4}\rfloor$. Since each time the problem is partitioned in Step 2, the number of integer variables decreases by one, the total number of iterations of the algorithm is upper bounded by
$\left(14(n+1)p2^{p(p-1)/4}\right)^{p}.$
Step 4: Termination. If the algorithm did not find any feasible solution to Equation MIQP, the algorithm correctly returns that the feasible region is empty. Otherwise, among all the feasible solutions to Equation MIQP found, the algorithm returns one with the minimum objective value. The result then follows from Lemma 15. This is because the function $f(x):=x^{\mathsf{T}}Hx+h^{\mathsf{T}}x$, from the original feasible region $\{x\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}:Wx\leq w\}$ to $\mathbb{R}$, is lower bounded by assumption, and thus has a (global) minimum, due to theorem 4 in [9]. ∎
@@
Acknowledgments: The author would like to thank two anonymous referees for comments and suggestions that improved the quality of this manuscript.